Heat is allowed to flow out of an ideal gas at constant volume so that it's pressure drops from 2.2atm to 1.4atm. Then the gas is expanded at constant pressure from a volume 6.8L to 9.3L until the temperature reaches it's final value.what is the total work done by the gas in the process,the internal energy of the gas and the total flow of heat into or out of the gas.
1
Expert's answer
2020-12-17T05:27:53-0500

The total work is done by the gas in the process

"W = P(V_2-V_1) \\

W = (1.4 \;atm)(1.01 \times 10^5 \;Pa)(9.3-6.8)(10^{-3} \;m^3/L) \\

W= 353.5 \;J"

The internal energy of the gas

Since there is no overall change in temperature there no change in internal energy.

The total flow of heat into or out of the gas

"Q = ΔU + W \\

Q = 0 + 353.5 \;J \\

Q = 353.5 \;J"